CWIS Queue Operations Documentation

Bruce Lewis November 22, 1995

Contents

1 Functional Overview 3

2 Server Characteristics 3 2.1 Compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Security . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 CPU Loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Network Loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 Requirements and Dependencies 3 3.1 Required . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2 Optional . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.3 Not Required . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4 Dependents 4

5 Building 4

6 Installation 4

7 Maintenance 5 7.1 Periodic Maintenance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 7.2 Relocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 7.3 Reconfiguration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 7.4 Restarts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 7.5 Updates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7.6 Shutdowns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7.7 Removal from Service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7.8 Backups and Restores . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

8 Troubleshooting 6 8.1 Diagnostics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 8.2 Known/Predicted Failure Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 8.3 Service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

9 State Information 6

1 Functional Overview

The CWIS queue is just a shared mh folder in AFS. When people on the cwis-request mailing list wants to answer questions, they scan that folder, refile a question or two into a private folder, and then answer. This prevents multiple people from answering at the same time.

The directory for this mh folder has permissions system:anyuser lik so that anybody can inc into the directory, but random people can’t read, erase or overwrite messages.

The cwis-help queue is exactly like the CWIS queue; just a different address and a different directory to inc into.

2 Server Characteristics

This service really is nothing more than a mail file on the server that gets inc’ed into the appropriate directory every 15 minutes or so. It makes sense to run this service on a server that

is already mkserv mail.

2.1 Compatibility

This service should not interfere with any other service.

2.2 Security

This service is as secure as feeding mail into discuss meetings.

2.3 CPU Loading

This is a low-load service. Its presence will be hardly noticeable.

2.4 Network Loading

No significant network load is introduced by this service, provided nobody is spamming the email address.

3 Requirements and Dependencies

This service depends on the existence of two directories, currently /afs/athena.mit.edu/org/c/cwis/Mail/inbox

and /afs/athena.mit.edu/org/c/cwis/Mail/cwis-help, with AFS permissions system:anyuser

lik. It relies on /usr/athena/bin/inc being available on the server, although if it’s sometimes unavailable, it will only delay things, not cause a mail bounce. No other mh library files or other executables are required.

3.1 Required

• srvtab - NO

• Kerberos - NO

• Athena - any release

• Platforms - anywhere cron, mkserv mail and inc are available

• Hesiod - NO

3.2 Optional

• syslog - sendmail will syslog any errors

3.3 Not Required

• root access - Does not need to run as root.

• name service - Does not depend on name services.

• disk space - Has no notable disk space requirements.

• swap space/real memory - Has no notable memory requirements.

• peripherals - requires no special peripherals

• /srvd - /usr/athena/bin/inc must be there

4 Dependents

If this service is unavailable, cwis-request and cwis-help people need to fall back on answering questions from the discuss meeting and cc’ing the group. No other services depend on this service.

5 Building

No code needs to be compiled.

6 Installation

1. Create a directory /Mail mode 770, group daemon. No mail is stored in this directory, but inc will complain if it can’t write /Mail/context.

2. Create one non-login passwd entry named “cwisq”, and another named “cwishelp”. The UID may be 1 or anything you like.

3. Add crontab entries like the following, except all on one line:

     0,15,30,45 * * * * /usr/athena/bin/inc
                           +/afs/athena.mit.edu/org/c/cwis/Mail/inbox
                           -file /usr/spool/mail/cwisq -truncate
     0,15,30,45 * * * * /usr/athena/bin/inc
                           +/afs/athena.mit.edu/org/c/cwis/Mail/cwis-help
                           -file /usr/spool/mail/cwishelp -truncate

They can be run as root or as cwisq/cwishelp, whichever you prefer.

4. Create a .private.sync with the following:

     copy usr/athena/bin -f
     copy usr/athena/bin/inc -p -f

so that inc is local.

7 Maintenance

7.1 Periodic Maintenance

Same as any other service that receives mail.

7.2 Relocation

Notify cwis-admin@mit.edu of the date of any relocation. If you are moving the service from oldhost to newhost, instruct them to type

   blanche cwis-request -a cwisq@oldhost.local -d cwisq@newhost.local
   blanche cwis-help -a cwishelp@oldhost.local -d cwishelp@newhost.local

7.3 Reconfiguration

If cwis-admin for some reason relocates the mh folder for the cwis queue, edit the crontab entry

file and replace /afs/athena.mit.edu/org/c/cwis/Mail/inbox with the new path. Similar for

cwis-help.

7.4 Restarts

In the unlikely event that sendmail dies, you may restart it with the following command:

/usr/lib/sendmail -bd -q30m

7.5 Updates

If an update is needed, follow the instructions for installation.

7.6 Shutdowns

No special shutdown procedure. Just remember that for as long as the service is down, mail piles up on the mailhub. It shouldn’t make a big difference for outages of less than a day. After 3 days mail starts bouncing.

7.7 Removal from Service

Give cwis-admin@mit.edu at least two weeks notice if you’re going to punt this service, and provide instructions for setting up the service on a private workstation.

7.8 Backups and Restores

No backups needed.

8 Troubleshooting

8.1 Diagnostics

If the service fails to work, edit the crontab entry to direct errors somewhere you can see them.

8.2 Known/Predicted Failure Modes

If /usr/spool/mail/cwisq does not exist, sendmail will not create it. Same for mail/cwishelp.

8.3 Service

Send bug reports to the bugs@mit.edu list.

9 State Information

The only local state information is that the last 15 minutes or less of mail is kept in /usr/spool/mail.